### v5

@@ 3.3 The $3\times 3$ Model
Let us now turn to the $3\times 3$ model of Example 2, where $\mathcal{Z}^{\ast}=\mathcal{T}^{\ast}=\{1,2\}$ and $\mathcal{Z}=\mathcal{T}=\{0,1,2\}$. We assume strict one-to-one targeting: for all of our results in this section, we impose Assumptions 1 - 5; $z=1$ targets $t=1$ and $z=2$ targets $t=2$.
The set $A$ in Corollary 1 can be ,$\emptyset,\{1\},\{2\}$, or $\{1,2\}$, with corresponding values of $t$ in $\{0\},\{0,1\},\{0,2\}$ or $\{0,1,2\}$ respectively. The set $c(\emptyset,0)$ corresponds to the never-takers $C_{000}$. For $A=\{1\}$ we get $C_{010}$ and $C_{111}$, and for $A=\{2\}$ we get $C_{002}$ and $C_{222}$. Finally, with $A=\{1,2\}$ we have $C_{012},C_{112}$, and $C_{212}$.
[…]
@@
Table 4: Response Groups of Example 2
@@
Unlike the $2\times 3$ model, even under strict one-to-one targeting the $3\times 3$ model does not satisfy unordered monotonicity. One could show it with the matrix algebra in Heckman and Pinto (2018). [^21] See Appendix D for details. It is more straightforward to note that when thee instrument value changes from $z=1$ to $z=2$, observations in $C_{010}$ move to treatment value 0, while observations in $C_{002}$ leave treatment 0. This is the definition of a two-way flow, which violates unordered monotonicity. Since the $3\times 3$ model has three instrument values and only two targeted treatments, Bai, Huang, Moon, Shaikh, and Vytlacil (2024, Example 4.7) shows that it satisfies their weaker general monotonicity assumption. As a consequence, the average potential outcomes $\mathbb{E}[Y_{i}(d)]$ can only be restricted by identification at infinity arguments.
@@ 3.3.1 Identification in the $3\times 3$ Model
We know from the third row of Table 3 that one restriction is missing to point-identify the probabilities of all eight response-groups. The following proposition shows that the probabilities of four of the eight elemental groups are point-identified: two groups of always-takers, and two groups of compliers. The other four probabilities are constrained by three adding-up constraints.
Proposition 7 (Response-group probabilities in the $3\times 3$ model under strict, one-to-one targeting).
@@
As before, the model has the following testable implications: $P(1|1)\geq P(1|0)\geq P(1|2)$, $P(2|2)\geq P(2|0)\geq P(2|1)$, and $P(0|0)\geq\max(P(0|1),P(0|2))$. The following proposition identifies a number of group average outcomes [^22] Again, these could also be derived using the general framework of Heckman and Pinto (2018), even though the unordered monotonicity assumption is not satisfied—see Appendix D..
Proposition 8 (Group average outcomes in the $3\times 3$ model under strict, one-to-one targeting).
@@
By itself, Proposition 8 does not allow us to identify an average treatment effect for any (even composite) response-group. Suppose for instance that we want to identify $\mathbb{E}(Y_{i}(1)-Y_{i}(0)|i\in C)$ for some group $C$. Then $C$ needs to exclude $C_{111}$, $C_{112}$, and $C_{212}$, since $\mathbb{E}(Y_{i}(0)|i\in C^{\prime})$ is not identified for any group $C^{\prime}$ that contains $C_{111}$, $C_{112}$, or $C_{212}$. Since we only know the mean outcome of treatment 1 for groups that contain one of these three elemental groups, the conclusion follows.
@@ 3.3.2 Using Positive Selection
Note that if we assumed $\mathbb{E}(Y_{i}(1)|i\in C_{112})=\mathbb{E}(Y_{i}(1)|i\in C_{212})$, then we could combine the two equations in the fourth displayed line of Proposition 8 and the probabilities of $C_{112}$ and $C_{212}$ (which are point-identified by Proposition 7) to obtain $\mathbb{E}(Y_{i}(1)|i\in C_{010}\operatorname*{{\bigcup}}C_{012})$. This would point-identify the average effect of treatment 1 vs treatment 0 on this composite complier group $C_{01\ast}$. While this assumption may be overly strong, it seems natural to impose that $Y_{i}(\tau)$ is on average larger in a response group that has $t=\tau$ for more values of $z$. Assumption 6 formalizes this intuition in our setting.
@@
Assumption 6 states a form of positive selection into treatment, as defined in Section 3.1. Consider Equation 3.8 for instance. It says that within the group of “1212 1212-compliers” $C_{\ast 12}=C_{012}\operatorname*{{\bigcup}}C_{112}\operatorname*{{\bigcup}}C_{212}$, those observations with $T(0)=1$ have a larger average counterfactual $Y(1)$ than those with $T(0)=2$. Corollary 3 shows that this gives bounds on the local average treatment effects for $C_{01\ast}$-compliers, with a similar result for Equation 3.9 and $C_{0\ast 2}$-compliers.
Corollary 3 (Identifying treatment effects in the $3\times 3$ model).
@@
    In both 1 and 2, “at least as large” can be replaced with “equals” if the corresponding inequality in Assumption 6 is an equality.
@@ 3.3.3 When is Positive Selection Plausible?
Now suppose that
Assumption 7 (Positive Codependence).
  1. (1)
    If $A$ and $B$ are two measurable sets in the $(\zeta,\xi)$ plane such that
    $(\zeta,\xi)\in A\text{ and }(\zeta^{\prime},\xi^{\prime})\in B\implies\left(\zeta\leq\zeta^{\prime}\text{ and }\xi^{\prime}\leq\xi^{\prime}\right),$
    then $\mathbb{E}(Y_{i}(2)|(\zeta_{i},\xi_{i})\in A)\leq\mathbb{E}(Y_{i}(2)|(\zeta_{i},\xi_{i})\in B)$.
  2. (2)
    If $A$ and $B$ are two measurable sets on the real line such that
    $\zeta\in A\text{ and }\zeta^{\prime}\in B\implies\zeta\leq\zeta^{\prime}$
    then $\mathbb{E}(Y_{i}(2)|\zeta_{i}\in A,\xi_{i}\geq b)\leq\mathbb{E}(Y_{i}(2)|\zeta_{i}\in B,\xi_{i}\geq b)$ for all $b$.
As shown in Figure 5, every point in $C_{112}^{(i)}$ has lower values of both $\zeta_{i}$ and $\xi_{i}$ than any point in $C_{212}$. Therefore, by Assumption 7(1), the expected value of $Y_{i}(2)$ in this triangle is smaller than $\mathbb{E}(Y_{i}(2)|i\in C_{212})$. Every point in $C_{112}^{(ii)}$ has is a smaller value of $\zeta_{i}$ than at any point in $C_{212}$ (fixing the value of $\xi_{i}\geq-\underline{U}(2)$ on both sides). Assumption 7(2) implies that the expected value of $Y_{i}(2)$ in this rectangle again is smaller than $\mathbb{E}(Y_{i}(2)|i\in C_{212})$. Combining these two inequalities gives $\mathbb{E}(Y_{i}(2)|i\in C_{212})\geq\mathbb{E}(Y_{i}(2)|i\in C_{112})$, that is (3.9).
Assumption 7 seems weak. Because both $\zeta_{i}=u_{i2}-u_{i1}$ and $\xi_{i}=u_{i2}-u_{i0}$ are increasing in $u_{i2}$, this assumption aligns well with the concept of positive selection. Suppose for instance that
$\mathbb{E}(Y_{i}(2)|u_{i0},u_{i1},u_{i2})-\mathbb{E}Y_{i}(2)=a_{0}u_{i0}+a_{1}u_{i1}+a_{2}u_{i2},$
where $a_{0}$, $a_{1}$, and $a_{2}$ are some constants, and that $(u_{i0},u_{i1},u_{i2})$ are jointly normal and mutually uncorrelated with the common mean 0 and the common variance 1. We interpret $(u_{i0},u_{i1},u_{i2})$ as the underlying primitive random variables that are normalized to have mean 0 and variance 1. It is easy to derive [^23] See Appendix G for details. that
$\displaystyle\mathbb{E}(Y_{i}(2)|\zeta_{i},\xi_{i})-\mathbb{E}Y_{i}(2)$ $\displaystyle=\frac{a_{2}+a_{0}-2a_{1}}{3}\zeta_{i}+\frac{a_{2}+a_{1}-2a_{0}}{3}\xi_{i}.$
Hence, in this example, Assumption 7 holds if and only if $a_{2}+a_{0}\geq 2a_{1}\text{ and }a_{2}+a_{1}\geq 2a_{0}.$ In summary, a sufficiently large value of $a_{2}$ induces positive selection, generating patterns similar to those of comparative advantage in generalized Roy models.
@@ 3.4.1 Monotonicity and Irrelevance
While we take quite a different path, our strict one-to-one targeting assumption turns out to yield exactly the same identifying restrictions as the combination of monotonicity and irrelevance in KLM. We show it in Appendix C.
@@
Proposition 9 (TSLS in the 3 $\times$ 3 model under strict, one-to-one targeting).
@@
Proposition 9 implies that $\beta_{1}$ and $\beta_{2}$ are weighted averages of the four local average treatment effects on the right-hand side of this system of two equations. The weights are functions of the four probabilities on the left-hand side, which are point identified by Proposition 7. However, these weights may be positive or negative. This complicates interpretation further [^24] Mogstad, Torgovitsky, and Walters (2021) give a set of assumptions under which the weights are positive in a model with multiple binary instruments..
@@ 3.4.2 Additional Assumptions
Using the additional information on next-best alternatives in KLM amounts, in our notation, to dropping the “intermediate” response-groups $C_{212}$ and $C_{112}$ from the data. Then the system of equations in Proposition 9 becomes diagonal and it yields
@@
Additional information of the type used by KLM often is not available. Moreover, assuming away $C_{112}$ and $C_{212}$ seems rather strong. On the other hand, reasonable assumptions can be used to generate bounds on the local average treatment effects for 1-compliers and 2-compliers. Corollary 4 illustrates this.
Corollary 4 (TSLS in the $3\times 3$ model under strict, one-to-one targeting).
@@
The regularity condition (3.10) ensures that the $2\times 2$ matrix that premultiplies $(\beta_{1},\beta_{2})^{\prime}$ in Proposition 9 be invertible262626 It holds if $C_{212}$ and $C_{112}$ have positive probability and either $C_{010}\operatorname*{{\bigcup}}C_{012}$ or $C_{002}\operatorname*{{\bigcup}}C_{012}$ has positive probability.. To interpret the assumptions on signs, suppose that $\mathcal{D}_{1}$ is positive. Since $\mathcal{C}_{1}=C_{010}\operatorname*{{\bigcup}}C_{012}\operatorname*{{\bigcup}}C_{212}$, the positivity of $\mathcal{D}_{1}$ states that the average effect of treatment 1 on $C_{010}\operatorname*{{\bigcup}}C_{012}\operatorname*{{\bigcup}}C_{212}$ is at least as large as on $C_{112}$. This is a form of positive selection that is in the same spirit as (but different from) Assumption 6. If this form of positive selection holds for both treatments, then the TSLS estimates overestimate the LATEs on the corresponding compliers if $\mathcal{D}>0$, and they underestimate them if $\mathcal{D}<0$.
To summarize, the TSLS estimators in the $3\times 3$ model are difficult to interpret unless additional information is available and/or some additional assumptions are imposed. If the groups $C_{112}$ and $C_{212}$ are indeed empty, then both the TSLS estimators and those we obtained in Corollary 3 should identify the LATEs on the 1- and 2-compliers. Comparing their values is a useful (if informal) way of testing the assumptions and of exploring further the heterogeneity of the treatment effects.
@@ 4 Empirical Example: The Head Start Impact Study
We now reexamine the Kline and Walters’s (2016) analysis of the Head Start Impact Study (HSIS) using our framework. We use exactly the same data as they did; we only apply different identifying assumptions.
Head Start is a federal program in the US that addresses various factors affecting children’s development in low-income families. It provides early childhood education (hereafter “preschool”) and health and nutrition services. HSIS was a longitudinal study conducted from 2002 to 2010 to assess the program’s impact on cognitive, social-emotional, and health outcomes. It focused on 84 communities where the demand for Head Start services was larger than the supply. HSIS randomly assigned about 5,000 three and four year old preschool children to either a treatment group which was offered Head Start services, or a control group which received no such offer. Children in either group could also attend other preschool centers if offered a slot
The structure of HSIS is identical to that of Example 1: it is a $2\times 3$ model. The treatments here consist of no preschool ($n$), Head Start ($h$), and other preschool centers ($c$): $\mathcal{T}=\{n,h,c\}$. The instrument is binary, with a control group ($z=0$) and a group that is offered admission to Head Start ($z=1$). The outcome variable is test scores, measured in standard deviations from their mean.
In the terminology of this paper, treatment assignment satisfies strict, one-to-one targeting: strict targeting as the instrument is binary, and one-to-one targeting as $z=1$ only targets Head Start [^27] Kamat (2024) analyzes HSIS using a different approach that focuses on how the choice sets available to a child vary with the value of the instrument.. Figure 6 reproduces Figure 3 in this setting. In addition to the three always-taker groups $C_{nn}$, $C_{cc}$, and $C_{hh}$, there are two complier groups: $C_{nh}$, and $C_{ch}$. In Sections 4.1 and 4.2, we focus on the LATEs on the two complier groups $C_{nh}$ and $C_{ch}$. Section 4.3 embeds the model into a larger, $3\times 3$ model in order to evaluate the marginal value of the public funds used in Head Start.
[…]

### v6

@@ 3.3 The $3\times 3$ Model
Let us now turn to the $3\times 3$ model of Example 2, where $\mathcal{Z}^{\ast}=\mathcal{T}^{\ast}=\{1,2\}$ and $\mathcal{Z}=\mathcal{T}=\{0,1,2\}$. We assume universal targeting: for all of our results in this section, we impose Assumptions 1, 2, 4, and 5; $z=1$ targets $t=1$ and $z=2$ targets $t=2$.
The set $A$ in Proposition˜3 can be $\emptyset,\{1\},\{2\}$, or $\{1,2\}$, with corresponding values of $t$ in $\{0\},\{0,1\},\{0,2\}$ or $\{0,1,2\}$ respectively. The set $c(\emptyset,0)$ corresponds to the never-takers $C_{000}$. For $A=\{1\}$ we get $C_{010}$ and $C_{111}$, and for $A=\{2\}$ we get $C_{002}$ and $C_{222}$. Finally, with $A=\{1,2\}$ we have $C_{012},C_{112}$, and $C_{212}$.
$u_{i1}-u_{i0}$$u_{i2}-u_{i0}$$C_{012}$$C_{010}$$C_{002}$$C_{000}$$C_{111}$$C_{222}$$C_{112}$$C_{212}$$P_{0}$$P_{1}$$P_{2}$
@@
Table 5: Response Groups of Example˜2
@@
Unlike the $2\times 3$ model, even under strict one-to-one targeting the $3\times 3$ model does not satisfy unordered monotonicity. One could show it with the matrix algebra in Heckman and Pinto (2018). [^20] See Appendix E for details. It is more straightforward to note that when thee instrument value changes from $z=1$ to $z=2$, observations in $C_{010}$ move to treatment value 0, while observations in $C_{002}$ leave treatment 0. This is the definition of a two-way flow, which violates unordered monotonicity. Since the $3\times 3$ model has three instrument values and only two targeted treatments, Bai, Huang, Moon, Shaikh, and Vytlacil (2024, Example 4.7) shows that it satisfies their weaker general monotonicity assumption. As a consequence, the average potential outcomes $\mathbb{E}[Y_{i}(d)]$ can only be restricted by identification at infinity arguments.
@@ 3.3.1 Identification in the $3\times 3$ Model
We know that one restriction is missing to point-identify the probabilities of all eight response-groups. The following proposition shows that the probabilities of four of the eight elemental groups are point-identified: two groups of always-takers, and two groups of compliers. The other four probabilities are constrained by three adding-up constraints.
Proposition 8 (Response-group probabilities in the $3\times 3$ model under universal targeting).
@@
As before, the model has the following testable implications: $P(1|1)\geq P(1|0)\geq P(1|2)$, $P(2|2)\geq P(2|0)\geq P(2|1)$, and $P(0|0)\geq\max(P(0|1),P(0|2))$. It follows from Bai, Huang, and Tabord-Meehan (2025, Example 3.3) that these implications are sharp.
The following proposition identifies a number of group average outcomes [^21] Again, these could also be derived using the general framework of Heckman and Pinto (2018), even though the unordered monotonicity assumption is not satisfied—see Appendix E..
Proposition 9 (Group average outcomes in the $3\times 3$ model under universal targeting).
@@
By itself, Proposition˜9 does not allow us to identify an average treatment effect for any (even composite) response-group. Suppose for instance that we want to identify $\mathbb{E}(Y_{i}(1)-Y_{i}(0)|i\in C)$ for some group $C$. Then $C$ needs to exclude $C_{111}$, $C_{112}$, and $C_{212}$, since $\mathbb{E}(Y_{i}(0)|i\in C^{\prime})$ is not identified for any group $C^{\prime}$ that contains $C_{111}$, $C_{112}$, or $C_{212}$. Since we only know the mean outcome of treatment 1 for groups that contain one of these three elemental groups, the conclusion follows.
@@ 3.3.2 Using Positive Selection
Note that if we assumed $\mathbb{E}(Y_{i}(1)|i\in C_{112})=\mathbb{E}(Y_{i}(1)|i\in C_{212})$, then we could combine the two equations in the fourth displayed line of Proposition˜9 and the probabilities of $C_{112}$ and $C_{212}$ (which are point-identified by Proposition˜8) to obtain $\mathbb{E}(Y_{i}(1)|i\in C_{010}\operatorname*{{\bigcup}}C_{012})$. This would point-identify the average effect of treatment 1 vs treatment 0 on this composite complier group $C_{01\ast}$. While this assumption may be overly strong, it seems natural to impose that $Y_{i}(\tau)$ is on average larger in a response group that has $t=\tau$ for more values of $z$. Assumption 6 formalizes this intuition in our setting.
@@
Assumption˜6 states a form of positive selection into treatment, as defined in Section 3.1. Consider Equation˜3.9 for instance. It says that within the group of $12$-compliers” $C_{\ast 12}=C_{012}\operatorname*{{\bigcup}}C_{112}\operatorname*{{\bigcup}}C_{212}$, those observations with $T(0)=1$ have a larger average counterfactual $Y(1)$ than those with $T(0)=2$. Proposition˜10 shows that this gives bounds on the local average treatment effects for $C_{01\ast}$-compliers, with a similar result for Equation˜3.10 and $C_{0\ast 2}$-compliers.
Proposition 10 (Identifying treatment effects in the $3\times 3$ model).
@@
    In both 1 and 2, “at least as large” can be replaced with “equals” if the corresponding inequality in Assumption˜6 is an equality.
The lower bounds given in Proposition˜10 are sharp relative to the restrictions imposed by the mean-independence of the instruments and the positive selection assumptions. Just as with the bounds derived in (3.8), this explicit sharp characterization obtains because the composite response-group probabilities
$\displaystyle\Pr(i\in C_{01\ast})$ $\displaystyle=\Pr(i\in C_{010}\operatorname*{{\bigcup}}C_{012})=P(0\mid 0)-P(0\mid 1),$
$\displaystyle\Pr(i\in C_{0\ast 2})$ $\displaystyle=\Pr(i\in C_{002}\operatorname*{{\bigcup}}C_{012})=P(0\mid 0)-P(0\mid 2),$
are point-identified, as established in Proposition˜8.
@@ 3.3.3 When is Positive Selection Plausible?
Suppose for instance that
$\mathbb{E}(Y_{i}(2)\mid u_{i0},u_{i1},u_{i2})-\mathbb{E}(Y_{i}(2))=a_{0}u_{i0}+a_{1}u_{i1}+a_{2}u_{i2},$
where $a_{0}$, $a_{1}$, and $a_{2}$ are constants, and $(u_{i0},u_{i1},u_{i2})$ are jointly normal and mutually uncorrelated with mean 0 and variance 1. Defining $\zeta_{i}=u_{i2}-u_{i1}$ and $\xi_{i}=u_{i2}-u_{i0}$, it is easy to see that
(3.11) $\displaystyle\mathbb{E}(Y_{i}(2)\mid\zeta_{i},\xi_{i})-\mathbb{E}(Y_{i}(2))$ $\displaystyle=\frac{a_{2}+a_{0}-2a_{1}}{3}\zeta_{i}+\frac{a_{2}+a_{1}-2a_{0}}{3}\xi_{i}.$
If $a_{2}\geq\max(a_{0},a_{1})+|a_{1}-a_{0}|$, then the coefficients of $\zeta_{i}$ and $\xi_{i}$ in Equation˜3.11 are both non-negative. It follows [^22] See Appendix H for details. from the geometry of Figure 5 that $\mathbb{E}(Y_{i}(2)\mid i\in C_{212})\geq\mathbb{E}(Y_{i}(2)\mid i\in C_{112})$. In summary, a sufficiently large value of $a_{2}$ induces positive selection, generating patterns similar to those of comparative advantage in generalized Roy models.
@@ 3.4.1 Monotonicity and Irrelevance
While we take quite a different path, our universal targeting assumption turns out to yield exactly the same identifying restrictions as the combination of monotonicity and irrelevance in KLM. We show it in Appendix˜D.
@@
Proposition 11 (TSLS in the 3 $\times$ 3 model under universal targeting).
@@
Proposition˜11 implies that $\beta_{1}$ and $\beta_{2}$ are weighted averages of the four local average treatment effects on the right-hand side of this system of two equations. The weights are functions of the four probabilities on the left-hand side, which are point identified by Proposition˜8. However, these weights may be positive or negative. This complicates interpretation further [^23] Mogstad, Torgovitsky, and Walters (2021) give a set of assumptions under which the weights are positive in a model with multiple binary instruments..
@@ 3.4.2 Additional Assumptions
Using the additional information on next-best alternatives in KLM amounts, in our notation, to dropping the “intermediate” response-groups $C_{212}$ and $C_{112}$ from the data. Then the system of equations in Proposition˜11 becomes diagonal and it yields
@@
Additional information of the type used by KLM often is not available. Moreover, assuming away $C_{112}$ and $C_{212}$ seems rather strong. On the other hand, reasonable assumptions can be used to generate bounds on the local average treatment effects for 1-compliers and 2-compliers. Proposition˜12 illustrates this.
Proposition 12 (TSLS in the $3\times 3$ model revisited).
@@
The regularity condition (3.12) ensures that the $2\times 2$ matrix that premultiplies $(\beta_{1},\beta_{2})^{\prime}$ in Proposition˜11 be invertible [^25] It holds if $C_{212}$ and $C_{112}$ have positive probability and either $C_{010}\operatorname*{{\bigcup}}C_{012}$ or $C_{002}\operatorname*{{\bigcup}}C_{012}$ has positive probability.. To interpret the assumptions on signs, suppose that $\mathcal{D}_{1}$ is positive. Since $\mathcal{C}_{1}=C_{010}\operatorname*{{\bigcup}}C_{012}\operatorname*{{\bigcup}}C_{212}$, the positivity of $\mathcal{D}_{1}$ states that the average effect of treatment 1 on $C_{010}\operatorname*{{\bigcup}}C_{012}\operatorname*{{\bigcup}}C_{212}$ is at least as large as on $C_{112}$. This is a form of positive selection that is in the same spirit as (but different from) Assumption˜6. If this form of positive selection holds for both treatments, then the TSLS estimates overestimate the LATEs on the corresponding compliers if $\mathcal{D}>0$, and they underestimate them if $\mathcal{D}<0$.
To summarize, the TSLS estimators in the $3\times 3$ model are difficult to interpret unless additional information is available and/or some additional assumptions are imposed. If the groups $C_{112}$ and $C_{212}$ are indeed empty, then both the TSLS estimators and those we obtained in Proposition˜10 should identify the LATEs on the 1- and 2-compliers. Comparing their values is a useful (if informal) way of testing the assumptions and of exploring further the heterogeneity of the treatment effects.
@@ 4 Empirical Example: The Head Start Impact Study
We now reexamine the Kline and Walters’s (2016) analysis of the Head Start Impact Study (HSIS) using our framework. We use exactly the same data as they did; we only apply different identifying assumptions [^26] Kamat (2024), Kamat, Norris, and Pecenco (2026) and Angrist, Santos, and Tecchio (2025) analyze the HSIS using different identifying approaches..
Head Start is a federal program in the US that addresses various factors affecting children’s development in low-income families. It provides early childhood education (hereafter “preschool”) and health and nutrition services. HSIS was a longitudinal study conducted from 2002 to 2010 to assess the program’s impact on cognitive, social-emotional, and health outcomes. It focused on 84 communities where the demand for Head Start services was larger than the supply. HSIS randomly assigned about 5,000 three and four year old preschool children to either a treatment group which was offered Head Start services, or a control group which received no such offer. Children in either group could also attend other preschool centers if offered a slot.
The structure of HSIS is identical to that of Example˜1: it is a $2\times 3$ model. The treatments here consist of no preschool ($n$), Head Start ($h$), and other preschool centers ($c$): $\mathcal{T}=\{n,h,c\}$. The instrument is binary, with a control group ($z=0$) and a group that is offered admission to Head Start ($z=1$). The outcome variable is test scores, measured in standard deviations from their mean.
In our notation, $\mathcal{Z}=\{0,1\}$ and $\mathcal{Z}^{\ast}=\{1\}$; the instrument $z=1$ uniquely and strictly targets the Head Start treatment $h$. In the terminology of this paper, this assignment structure satisfies universal targeting. The set of inequalities in (2.4) is empty, as $z=0$ is the only non-targeting instrument value. On the other hand, (2.5) yields the following sharp testable implications:
$P(n|1)<P(n|0)\quad\text{and}\quad P(c|1)<P(c|0).$
In words, the Head Start offer must strictly reduce the choice probabilities of the other two alternatives (no preschool $n$ and other centers $c$). The data is fully consistent with these implications, as evidenced by the proportions of the two complier groups reported in Panel A of Table˜6 (see also Kline and Walters (2016, Table 3)).
Figure 6 reproduces Figure 3 in this setting. In addition to the three always-taker groups $C_{nn}$, $C_{cc}$, and $C_{hh}$, there are two complier groups: $C_{nh}$, and $C_{ch}$. In Sections 4.1 and 4.2, we focus on the LATEs on the two complier groups $C_{nh}$ and $C_{ch}$. Section 4.3 embeds the model into a larger, $3\times 3$ model in order to evaluate the marginal value of the public funds used in Head Start.
$u_{ih}-u_{in}$$u_{ic}-u_{in}$$C_{nh}$$C_{nn}$$C_{cc}$$C_{hh}$$C_{ch}$$P_{n}$$P_{h}$